OpenStudy (anonymous):

systems of equations, simple elimination method the sum of two numbers is 19 and ther difference is 9. what are the numbers? and how do you solve it by using the elimination method??

OpenStudy (anonymous):

x + y = 19 x - y = 9 add the 2 equations to eliminate y: 2x = 28 x = 14 plug x=14 in first equation: 14 + y = 19 y = 19-14 = 5 2 numbers are 5 and 14

OpenStudy (anonymous):

You want to eliminate a variable. Sometimes you have to multiply an equation by a negative number, but not in this case. Simply add the 2 equations together to eliminate y. x+y=19 x-y=9 2x=28 x=14 14+y=19 y=5 Check: 14-5=9, true